OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5)with a slope of -3. Answer y = -3x - 13 y = -3x + 11 y = -3x + 1 y = -3x + 13

OpenStudy (lgbasallote):

use the point slope form y- y1 = m(x-x1) y1 = -5 x1 = 2 m = -3 substitute then turn to general form

OpenStudy (radar):

Just plug 2 for x and -5 for y and see which equation is true.

OpenStudy (mani_jha):

Or, since the slopes are -3 in all the options, you can just substitute x=2 and y=-5 and check which option is correct.